### v1

@@ Methods
For a subset of edges $N\subset E$ representing the set of faulty locations, we define the set of level-$k$ clusters $\mathsf{N}_{k}$ using the monotonically increasing sequences $\quantity{b_{k}}_{k=1,2,\ldots},\quantity{d_{k}}_{k=1,2,\ldots}\subset\mathbb{R}_{\geq 0}$, as shown in Definition S5 in the Supplementary Information. Then, we show the following lemma.
@@
Let $G=(V,E)$ be a graph, and assume that a set $N$ of faulty locations is randomly specified as a subset of the edges of $G$ in such a way that there exists a rate $\tilde{p}$ such that any subset $S$ of edges satisfies $|\mathbb{P}\quantity[N\supset S]\leq\tilde{p}^{\absolutevalue{S}}$. Suppose that there exists constants $\Lambda,\Delta>0$ such that
@@
$p_{k}\leq\tilde{p}^{2^{k}}\prod_{j=0}^{k-1}\left[\Lambda\quantity(b_{k-j}+\frac{d_{k-j}}{2})^{\Delta}\right]^{2^{j}}.$ (8)
@@
Lemma 6 implies that if $p$ in Eq. (4) is below a nonzero threshold $p_{\mathrm{th}}>0$, then $p_{k}$ decreases doubly exponentially with $k$ for some choice of $b_{k},d_{k}$, as shown by the following lemma. Note that the value $p_{\rm th}$ given here is a lower bound on the true threshold, and such analytical lower bounds are usually not tight.
@@
using a scale parameter $\lambda>1$, positive constants $\beta,\gamma>0$, and a monotonically increasing function $f:\mathbb{Z}_{>0}\to\mathbb{R}_{>0}$ satisfying
$\sum_{j=0}^{k-1}f(k-j+1)2^{j}=c\cdot 2^{k}+o(2^{k}),$ (10)
@@
$p_{k}$ $\leq\tilde{p}^{2^{k}}\prod_{j=0}^{k-1}\left[\Lambda\left(\beta\lambda^{f(k-j+1)}+{\gamma\lambda^{f(k-j)}+1\over 2}\right)^{\Delta}\right]^{2^{j}}$ (13)
$\leq\quantity(\xi p)^{2^{k}}\prod_{j=0}^{k-1}\left[\Lambda\left(\beta+{\gamma+\lambda^{-f(1)}\over 2}\right)^{\Delta}\lambda^{\Delta f(k-j+1)}\right]^{2^{j}}$ (14)
@@
$=O\quantity((p/p_{\mathrm{th}})^{2^{k}}).$ (16)
@@
We define the sequence $\quantity{f_{k}}_{k=1,2,\ldots}$ by
$f_{k}$ $\coloneqq\begin{cases}1&(k=1)\\ 1-\sum_{k^{\prime\prime}=1}^{k-1}{(f_{k^{\prime\prime}}+1)(d_{k^{\prime\prime}}+1)\over(f_{k^{\prime\prime}}+1)(d_{k^{\prime\prime}}+1/2)+f_{k^{\prime\prime}}(b_{k^{\prime\prime}}-1)}&(k\geq 2)\end{cases}.$ (19)
@@
We show this lemma by induction on $k$. For $k=1$, any level-$1$ cluster $C$ has diameter at most $d_{1}+1$, and buffer at least $b_{1}-1=f_{1}(b_{1}-1)>d_{1}+1$. Therefore, any level-$1$ extended UF cluster $\tilde{C}_{t}$ stops growing before merging with other clusters, and $m_{1}\leq(d_{1}+1)/2=(d_{1}+1)/2f_{1}$ holds.
For $k>1$, we assume the statement holds for all $k^{\prime}<k$. By assumption, when a level-$k$ extended UF cluster $\tilde{C}_{t}$ overlapping with a level-$k$ cluster $C$ of error configuration merges with a level-$k^{\prime\prime}$ ($k^{\prime\prime}<k$) extended UF cluster $\tilde{C}^{\prime\prime}_{t}$, the level-$k^{\prime\prime}$ extended UF cluster $\tilde{C}^{\prime\prime}_{t}$ is stable. Suppose that $\tilde{C}_{t}$ merges, at step $t$ before stopping the growth, with a level-$k^{\prime}$ ($k^{\prime}\geq k$) extended UF cluster $\tilde{C}^{\prime}_{t}$ overlapping with a level-$k^{\prime}$ cluster $C^{\prime}$ of error configuration. If $t>d_{k}+1$ holds, the level-$k$ extended UF cluster $\tilde{C}_{t}$ must be stable since all the level-$k$ clusters of error configuration in $\tilde{C}_{t}$ have diameter $d_{k}+1$ smaller than $t$ steps of growing UF clusters in $\tilde{C}^{\prime}_{t}$. Thus, for $\tilde{C}_{t}$ to keep growing, we necessarily have
@@
$1-\sum_{k^{\prime\prime}=1}^{k-1}{d_{k^{\prime\prime}}+1+2m_{k^{\prime\prime}}\over d_{k^{\prime\prime}}+1+m_{k^{\prime\prime}}+b_{k^{\prime\prime}}-1}$
$\geq 1-\sum_{k^{\prime\prime}=1}^{k-1}{d_{k^{\prime\prime}}+1+{d_{k^{\prime\prime}}+1\over f_{k^{\prime\prime}}}\over d_{k^{\prime\prime}}+1+{d_{k^{\prime\prime}}+1\over 2f_{k^{\prime\prime}}}+b_{k^{\prime\prime}}-1}$ (25)
@@
To exploit this stopping guarantee on the UF decoder together with the error clustering result in Lemma 7, we take the sequences $\{(b_{k},d_{k})\}_{k=1,2,\ldots}$ as
$b_{k}=\beta\lambda^{(k+1)\log(k+1)}+1,\quad d_{k}=\gamma\lambda^{k\log k}-1,$ (30)
which corresponds to $f(k)=k\log k$ in Eq. (9), satisfying Eq. (10) for $c=\sum_{n=2}^{\infty}{n\log n\over 2^{n-1}}\approx 3.57257$, where $\log$ is the logarithm to the base of Napier’s constant $e\coloneqq\lim_{n\to\infty}(1+{1\over n})^{n}\approx 2.71828$. We assume that the constants satisfy the condition that
$\lambda>e,\quad{8\gamma\over\beta}\left[\zeta(\log\lambda)-1\right]\leq 1,$ (31)
@@
Then, $f_{k}$ defined in Eq. (19) satisfies $f_{k}\geq{1\over 2}$, which is shown by induction on $k$ as follows. For $k=1$, $f_{1}=1\geq{1\over 2}$. For $k>1$, if $f_{k^{\prime}}\geq{1\over 2}$ holds for all $k^{\prime}<k$, we have
$f_{k}$ $\geq 1-\sum_{k^{\prime\prime}=1}^{k-1}{2(d_{k^{\prime\prime}}+1)\over f_{k^{\prime\prime}}(b_{k^{\prime\prime}}-1)}$ (32)
$\geq 1-{4\gamma\over\beta}\left(\sum_{k^{\prime\prime}=2}^{\infty}\lambda^{-\log k^{\prime\prime}}\right)$ (33)
$=1-{4\gamma\over\beta}\left[\zeta(\log\lambda)-1\right]$ (34)
@@
$\begin{split}\beta\lambda>2\gamma,\quad\gamma\geq{2\gamma\over\lambda}+2\beta+\lambda^{-2\log 2},\quad\gamma\lambda\geq 2.\end{split}$ (36)
Under the above choice of $\{(b_{k},d_{k})\}_{k\geq 1}$ and the constants, we show the threshold theorem.
@@
We estimate a lower bound on the threshold $p_{\mathrm{th}}$ for the UF decoder. We numerically search the parameters $\beta,\gamma,\lambda$ to maximize $p_{\mathrm{th}}$ under the constraints in Eqs. (6) and (31) to find a parameter choice of
$\beta\approx 1.2,\quad\gamma\approx 2.8,\quad\lambda\approx 107,$ (46)
and substituting these parameters in Eq. (12) yields
$p_{\mathrm{th}}\gtrapprox 2.5\times 10^{-26}/\xi,$ (47)
@@
As shown in the main text, level-kkextended UF clusters can be decoded in parallel in time $O(\operatorname{poly}(d_{k}))\leq\lambda^{\tau k\log k}$ for a sufficiently large constant $\tau$, where we use Eq. (30). The probability of having a level-$k$ extended UF cluster in the decoding window is upper-bounded by
@@
We take the sequences $\quantity{(b_{k},d_{k})}_{k=1,2,\ldots}$ as
@@
$\gamma\lambda\geq 2,\quad\beta\lambda>\gamma,\quad\gamma\lambda-\lambda^{-1}\geq 2(\gamma+2\beta\lambda).$ (62)
If
@@ Acknowledgments
EL and HY discussed part of this work during the KIAS workshop on Quantum Information Dynamics and Non-Equilibrium Physics (Quantum Information and Dynamics 2025). HY acknowledges Nicolas Delfosse and Shouzhen Gu for discussions at the $7$th International Conference on Quantum Error Correction (QEC25), and Naomi Nickerson for discussions at the Seeking Quantum Advantage Workshop and Conference (SEEQA2025). EL thanks Nathaniel Selub for discussions and collaboration on related work. SY was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number 23KJ0734, FoPM, WINGS Program, the University of Tokyo, and DAIKIN Fellowship Program, the University of Tokyo. EL was supported by a Miller Fellowship. HY was supported by JST PRESTO Grant Number JPMJPR201A, JPMJPR23FC, JSPS KAKENHI Grant Number JP23K19970, and JST CREST Grant Number JPMJCR25I5.
@@
Let $G=(V,E)$ be a graph and $N\subset E$ be a subset of edges. We consider monotonically increasing sequences $\{d_{k}\}_{k=1,2,\ldots}$ and $\{b_{k}\}_{k=1,2,\ldots}$ satisfying
@@
The level-$k$ clustered set $\mathsf{N}_{k}$ and level-$k$ isolated set $\mathcal{N}_{k}$ are defined inductively as
@@
We also define sets $\mathsf{E}_{k}$ and $\mathcal{E}_{k}$ by
@@
$N$ $\subset B_{e}(d_{k}/2)\quad\text{for all $N\in M_{k}(e)$},$ (S39)
@@
$|\displaystyle\absolutevalue{M_{k}(e)}$ $\leq\prod_{j=0}^{k-1}\left[\Lambda\quantity(b_{k-j}+\frac{d_{k-j}}{2})^{\Delta}\right]^{2^{j}}.$ (S42)
@@
We will first show Eq. (S39) by induction on $k$. Since $M_{1}(e)=\{\{e\}\}$, Eq. (S39) holds for $k=1$. Assuming that Eq. (S39) holds for $k$, we will show Eq. (S39) for $k+1$. If $e\in\mathcal{N}_{k+1}$ holds, then by definition, there exists $e^{\prime}\in\mathcal{N}_{k}$ such that $e$ and $e^{\prime}$ are $(d_{k}/2,b_{k}+d_{k}/2)$-linked. By definition of $M_{k+1}(e)$, any $N\in M_{k+1}(e)$ includes, as subsets, some $N_{1}\in M_{k}(e)$ and some $N_{2}\in M_{k}(e^{\prime})$ for an edge $e^{\prime}$ such that $e$ and $e^{\prime}$ are $(d_{k}/2,b_{k}+d_{k}/2)$-linked; moreover, by minimality, $N$ contains no elements other than those in $N_{1}$ and $N_{2}$. Thus, we obtain
@@
By assumption, we have $N_{1}\subset B_{e}(d_{k}/2)$ and $N_{2}\subset B_{e^{\prime}}(d_{k}/2)$. Therefore, we obtain
$N\subset B_{e}(d_{k}+b_{k})\subset B_{e}(d_{k+1}/2),$ (S44)
where the factor $d_{k}+b_{k}$ arises because $e$ and $e^{\prime}$ are $(d_{k}/2,b_{k}+d_{k}/2)$-linked.
We then show Eq. (S40) similarly to Eq. (S39) by induction on $k$. Since $M_{0}(e)=\{\{e\}\}$, Eq. (S40) holds for $k=0$. Assuming that Eq. (S40) holds for $k$, we will show Eq. (S40) for $k+1$. In the decomposition (S43), since $N_{1}\subset B_{e}(d_{k}/2)$, $N_{2}\subset B_{e^{\prime}}(d_{k}/2)$, and $\mathrm{dist}(e,e^{\prime})>d_{k}$ hold, we have $N_{1}\cap N_{2}=\varnothing$. Therefore, we obtain
@@
We finally show Eq. (S42). Since any $N\in M_{k}(e)$ is given by $N=N_{1}\cup N_{2}$ such that $N_{1}\in M_{k-1}(e)$ and $N_{2}\in M_{k-1}(e^{\prime})$, and $e$ and $e^{\prime}$ are $(d_{k}/2,b_{k}+d_{k}/2)$-linked, we obtain
@@
$|\displaystyle\absolutevalue{M_{k}(e)}$ $\leq\prod_{j=0}^{k-1}\left[\Lambda\quantity(b_{k-j}+\frac{d_{k-j}}{2})^{\Delta}\right]^{2^{j}}.$ (S50)

### v2

@@ Discussion
Extension of other codes
Our proof techniques for the threshold theorem for the UF decoder for the surface code, in principle, apply to quantum low-density parity-check (QLDPC) codes satisfying the appropriate spacetime locality condition. For example, it would be promising to use our framework to analyze various decoders for higher-dimensional surface or color codes. QLDPC codes that do not satisfy spacetime locality, such as quantum expander codes [60, 61], lifted-product codes [22, 62], bivariate bicycle codes [63], and asymptotically good QLDPC codes [64, 65, 66], do not directly satisfy our assumptions. Nevertheless, it would be interesting to identify alternative conditions under which the error-clustering framework can be applied to these codes in future work.
@@ Methods
Given a subset $N\subset E$ of edges representing faulty locations, we introduce level-$k$ clusters as follows (see Definition S5 in the Supplementary Information for the precise definition). For parameters $d,b>0$, a $(d,b)$-cluster in a subset $N\subset E$ of edges is a connected subset of $N$ in the graph with diameter at most $d$, surrounded by a width-$b$ buffer region that contains no edges of $N$. For level $k\in\{0,1,\ldots\}$, we choose increasing sequences of parameters $\quantity{d_{k}}_{k=0,1,\ldots}$ and $\quantity{b_{k}}_{k=0,1,\ldots}$. We then define the residual set $\mathsf{N}_{k}$ of unclustered edges inductively by $\mathsf{N}_{0}=N$, with $\mathsf{N}_{k+1}$ given by the set of edges in $\mathsf{N}_{k}$ that are not contained in any $(d_{k},b_{k})$-cluster in $\mathsf{N}_{k}$. The set $\mathsf{E}_{k}=\mathsf{N}_{k}\setminus\mathsf{N}_{k+1}$ consists of the edges belonging to level-$k$ clusters, with all edges in level-$k^{\prime}$ ($k^{\prime}<k$) clusters already removed. For a general choice of $d_{k}$ and $b_{k}$, the following lemma provides a high-probability bound on the occurrence of level-$k$ clusters.
@@
Let $G=(V,E)$ be a graph, and assume that a set $N$ of faulty locations is randomly specified as a subset of the edges of $G$ in such a way that there exists a rate $\tilde{p}$ such that any subset $S$ of edges satisfies $|\mathbb{P}\quantity[N\supset S]\leq\tilde{p}^{\absolutevalue{S}}$. Suppose that there exist constants $\Lambda,\Delta>0$ such that the locality condition
@@
$p_{k}\leq\tilde{p}^{2^{k}}\prod_{j=0}^{k-1}\left[\Lambda\quantity(b_{k-j-1}+\frac{d_{k-j-1}}{2})^{\Delta}\right]^{2^{j}}.$ (8)
@@
Lemma 6 is essential for a threshold theorem: if $p$ in Eq. (4) is below a nonzero threshold $p_{\mathrm{th}}>0$, then $p_{k}$ decreases doubly exponentially with $k$. The value $p_{\rm th}$ obtained in this way is a rigorous lower bound on the true threshold, although such analytical lower bounds are usually not tight. Existing threshold analyses [15, 14, 19, 20] take $b_{k}$ and $d_{k}$ to grow with the same scaling up to constant factors. By contrast, we show here that the threshold theorem still holds even when $b_{k}$ grows faster than $d_{k}$; more precisely, for a constant $\lambda$, we can take $b_{k}=\Theta(\lambda^{f(k+1)})$ and $d_{k}=\Theta(\lambda^{f(k)})$ even for an increasing function $f(k)\gg k$, so that the buffer width $b_{k}$, with exponent $f(k+1)$, can be substantially larger than the cluster diameter $d_{k}$, with exponent $f(k)$, with their relative separation controlled by the growth rate of $f$. For example, in our analysis of the UF decoder, we take $f(k)\approx k\log k$, while the framework itself permits even faster-growing choices such as $f(k)=k^{\alpha}$ for $\alpha>1$.
@@
using a scale parameter $\lambda>1$, positive constants $\beta,\gamma>0$, and a monotonically increasing function $f:\mathbb{Z}_{\geq 0}\to\mathbb{R}_{\geq 0}$ satisfying
$\sum_{j=0}^{k-1}f(k-j)2^{j}=c\cdot 2^{k}+o(2^{k}),$ (10)
@@
$p_{k}$ $\leq\tilde{p}^{2^{k}}\prod_{j=0}^{k-1}\left[\Lambda\left(\beta\lambda^{f(k-j)}+{\gamma\lambda^{f(k-j-1)}+1\over 2}\right)^{\Delta}\right]^{2^{j}}$ (13)
$\leq\quantity(\xi p)^{2^{k}}\prod_{j=0}^{k-1}\left[\Lambda\left(\beta+{\gamma+\lambda^{-f(1)}\over 2}\right)^{\Delta}\lambda^{\Delta f(k-j)}\right]^{2^{j}}$ (14)
@@
$=O\quantity((p/p_{\mathrm{th}})^{2^{k}}),$ (16)
where we use the monotonicity of the function $f$ given by $f(k-j)\geq f(k-j-1)$ and $f(1)\leq f(k-j)$ for $j\leq k-1$ in Eq. (14). ∎
@@
We define the sequence $\quantity{f_{k}}_{k=0,1,\ldots}$ by
$f_{k}$ $\coloneqq\begin{cases}1&(k=0)\\ 1-\sum_{k^{\prime\prime}=0}^{k-1}{(f_{k^{\prime\prime}}+1)(d_{k^{\prime\prime}}+1)\over(f_{k^{\prime\prime}}+1/2)(d_{k^{\prime\prime}}+1)+f_{k^{\prime\prime}}(b_{k^{\prime\prime}}-1)}&(k\geq 1)\end{cases}.$ (19)
@@
We show this lemma by induction on $k$. For $k=0$, any level-0 cluster $C$ has diameter at most $d_{0}+1$, and buffer at least $b_{0}-1=f_{0}(b_{0}-1)>d_{0}+1$. Therefore, any level-0 extended UF cluster $\tilde{C}_{t}$ stops growing before merging with other clusters, and $m_{0}\leq(d_{0}+1)/2=(d_{0}+1)/2f_{0}$ holds.
For $k\geq 1$, we assume the statement holds for all $k^{\prime}<k$. By assumption, when a level-$k$ extended UF cluster $\tilde{C}_{t}$ overlapping with a level-$k$ cluster $C$ of error configuration merges with a level-$k^{\prime\prime}$ ($k^{\prime\prime}<k$) extended UF cluster $\tilde{C}^{\prime\prime}_{t}$, the level-$k^{\prime\prime}$ extended UF cluster $\tilde{C}^{\prime\prime}_{t}$ is stable. Suppose that $\tilde{C}_{t}$ merges, at step $t$ before stopping the growth, with a level-$k^{\prime}$ ($k^{\prime}\geq k$) extended UF cluster $\tilde{C}^{\prime}_{t}$ overlapping with a level-$k^{\prime}$ cluster $C^{\prime}$ of error configuration. If $t>d_{k}+1$ holds, the level-$k$ extended UF cluster $\tilde{C}_{t}$ must be stable since all the level-$k$ clusters of error configuration in $\tilde{C}_{t}$ have diameter $d_{k}+1$ smaller than $t$ steps of growing UF clusters in $\tilde{C}^{\prime}_{t}$. Thus, for $\tilde{C}_{t}$ to keep growing, we necessarily have
@@
$1-\sum_{k^{\prime\prime}=0}^{k-1}{d_{k^{\prime\prime}}+1+2m_{k^{\prime\prime}}\over d_{k^{\prime\prime}}+1+m_{k^{\prime\prime}}+b_{k^{\prime\prime}}-1}$
$\geq 1-\sum_{k^{\prime\prime}=0}^{k-1}{d_{k^{\prime\prime}}+1+{d_{k^{\prime\prime}}+1\over f_{k^{\prime\prime}}}\over d_{k^{\prime\prime}}+1+{d_{k^{\prime\prime}}+1\over 2f_{k^{\prime\prime}}}+b_{k^{\prime\prime}}-1}$ (25)
@@
To exploit this stopping guarantee on the UF decoder together with the error clustering result in Lemma 7, we take the sequences $\{(b_{k},d_{k})\}_{k=0,1,\ldots}$ as
$b_{k}=\beta\lambda^{(k+2)\log(k+2)}+1,\quad d_{k}=\gamma\lambda^{(k+1)\log(k+1)}-1,$ (30)
which corresponds to $f(k)=(k+1)\log(k+1)$ in Eq. (9), satisfying Eq. (10) for $c=\sum_{n=2}^{\infty}{2n\log n\over 2^{n}}\approx 3.57256$, where $\log$ is the logarithm to the base of Napier’s constant $e\coloneqq\lim_{n\to\infty}(1+{1\over n})^{n}\approx 2.71828$. We assume that the constants satisfy the condition that
$\lambda>e,\quad{6\gamma\over\beta}[\zeta(\log\lambda)-1+\lambda^{-2\log 2}]\leq 1,$ (31)
@@
Then, $f_{k}$ defined in Eq. (19) satisfies $f_{k}\geq{1\over 2}$, which is shown by induction on $k$ as follows. For $k=0$, $f_{0}=1\geq{1\over 2}$ holds. For $k\geq 1$, if $f_{k^{\prime}}\geq{1\over 2}$ holds for all $k^{\prime}<k$, we have
$f_{k}$ $\geq 1-\sum_{k^{\prime\prime}=0}^{k-1}{{3\over 2}(d_{k^{\prime\prime}}+1)\over d_{k^{\prime\prime}}+1+{1\over 2}(b_{k^{\prime\prime}}-1)}$ (32)
$\geq 1-{3\gamma\over\beta}\sum_{k^{\prime\prime}=0}^{k-1}\lambda^{(k^{\prime\prime}+1)\log(k^{\prime\prime}+1)-(k^{\prime\prime}+2)\log(k^{\prime\prime}+2)}$ (33)
$\geq 1-{3\gamma\over\beta}\left(\lambda^{-2\log 2}+\sum_{k^{\prime\prime}=1}^{\infty}\lambda^{-\log(k^{\prime\prime}+1)}\right)$ (34)
$=1-{3\gamma\over\beta}[\zeta(\log\lambda)-1+\lambda^{-2\log 2}]$ (35)
@@
$2\lambda^{-2\log 2}\leq{\beta\over\gamma}\leq{1-4\lambda^{-2\log 2}\over 3},\quad\gamma\geq 2.$ (37)
Under the above choice of $\{(b_{k},d_{k})\}_{k=0,1,\ldots}$ and the constants, we show the threshold theorem.
@@
We estimate a lower bound on the threshold $p_{\mathrm{th}}$ for the UF decoder. We take the constants $\beta,\gamma,\lambda$ satisfying
${\beta\over\gamma}={1-4\lambda^{-2\log 2}\over 3},\quad\gamma=2,$ (47)
${1-4\lambda^{-2\log 2}\over 3}=6[\zeta(\log\lambda)-1+\lambda^{-2\log 2}]>2\lambda^{-2\log 2}.$ (48)
For reference, these values are
$\beta\approx 0.66,\quad\gamma=2,\quad\lambda\approx 93.$ (49)
Substituting these parameters in Eq. (12) yields
$p_{\mathrm{th}}>6.6\times 10^{-25}/\xi,$ (50)
@@
As shown in the main text, level-$k$ extended UF clusters can be decoded in parallel in time $O(\operatorname{poly}(d_{k}))\leq\lambda^{\tau k\log k}$ for a sufficiently large constant $\tau$, where we use Eq. (30). The probability of having a level-$k$ extended UF cluster in the decoding window is upper-bounded by
@@
We take the sequences $\quantity{(b_{k},d_{k})}_{k=0,1,\ldots}$ as
@@
$\lambda\geq 2,\quad\lambda^{-1}<{\beta\over\gamma}\leq{1-4\lambda^{-1}\over 3}.$ (65)
Note that the constants $\beta,\gamma,\lambda$ satisfying the above constraints exist since $\lambda^{-1}<{1-4\lambda^{-1}\over 3}$ holds for $\lambda>7$. If
@@
Data availability
No data is used in this study.
Code availability
No code is used in this study.
@@ Acknowledgments
EL and HY discussed part of this work during the KIAS workshop on Quantum Information Dynamics and Non-Equilibrium Physics (Quantum Information and Dynamics 2025). HY acknowledges Nicolas Delfosse and Shouzhen Gu for discussions at the $7$th International Conference on Quantum Error Correction (QEC25), and Naomi Nickerson for discussions at the Seeking Quantum Advantage Workshop and Conference (SEEQA2025). EL thanks Nathaniel Selub for discussions and collaboration on related work. We thank Ryo Mikami for pointing out typographical errors in the manuscript. SY was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number 23KJ0734, FoPM, WINGS Program, the University of Tokyo, and DAIKIN Fellowship Program, the University of Tokyo. EL was supported by a Miller Fellowship. HY was supported by JST PRESTO Grant Number JPMJPR201A, JPMJPR23FC, JSPS KAKENHI Grant Number JP23K19970, JST CREST Grant Number JPMJCR25I5, JST [Moonshot R&D] [Grant Number JPMJMS256J], and Faculty Research Funding from Google Quantum AI.
@@
Competing Interests
The authors declare no competing interests.
@@
Let $G=(V,E)$ be a graph and $N\subset E$ be a subset of edges. We consider monotonically increasing sequences $\{d_{k}\}_{k=0,1,\ldots}$ and $\{b_{k}\}_{k=0,1,\ldots}$ satisfying
@@
We inductively define the level-$k$ residual sets $\mathsf{N}_{k}$ and $\mathcal{N}_{k}$ of unclustered and non-isolated edges, respectively, as follows.
@@
Using these sets, we define the level-$k$ clustered edge set $\mathsf{E}_{k}$ and level-$k$ isolated edge set $\mathcal{E}_{k}$ by
@@
$N$ $\subset B_{e}(d_{k}/4)\quad\text{for all $N\in M_{k}(e)$},$ (S39)
@@
$|\displaystyle\absolutevalue{M_{k}(e)}$ $\leq\prod_{j=0}^{k-1}\left[\Lambda\quantity(b_{k-j-1}+\frac{d_{k-j-1}}{2})^{\Delta}\right]^{2^{j}}.$ (S42)
@@
We will first show Eq. (S39) by induction on $k$. Since $M_{0}(e)=\{\{e\}\}$, Eq. (S39) holds for $k=0$. Assuming that Eq. (S39) holds for $k$, we will show Eq. (S39) for $k+1$. If $e\in\mathcal{N}_{k+1}$ holds, then by definition, there exists $e^{\prime}\in\mathcal{N}_{k}$ such that $e$ and $e^{\prime}$ are $(d_{k}/2,b_{k}+d_{k}/2)$-linked. By definition of $M_{k+1}(e)$, any $N\in M_{k+1}(e)$ includes, as subsets, some $N_{1}\in M_{k}(e)$ and some $N_{2}\in M_{k}(e^{\prime})$ for an edge $e^{\prime}$ such that $e$ and $e^{\prime}$ are $(d_{k}/2,b_{k}+d_{k}/2)$-linked; moreover, by minimality, $N$ contains no elements other than those in $N_{1}$ and $N_{2}$. Thus, we obtain
@@
By assumption, we have $N_{1}\subset B_{e}(d_{k}/4)$ and $N_{2}\subset B_{e^{\prime}}(d_{k}/4)$. Therefore, we obtain
$N\subset B_{e}(3d_{k}/4+b_{k})\subset B_{e}(d_{k+1}/4),$ (S44)
where the factor $3d_{k}/4+b_{k}$ arises because $e$ and $e^{\prime}$ are $(d_{k}/2,b_{k}+d_{k}/2)$-linked.
We then show Eq. (S40) similarly to Eq. (S39) by induction on $k$. Since $M_{0}(e)=\{\{e\}\}$, Eq. (S40) holds for $k=0$. Assuming that Eq. (S40) holds for $k$, we will show Eq. (S40) for $k+1$. In the decomposition (S43), since $N_{1}\subset B_{e}(d_{k}/4)$, $N_{2}\subset B_{e^{\prime}}(d_{k}/4)$, and $\mathrm{dist}(e,e^{\prime})>d_{k}/2$ hold, we have $N_{1}\cap N_{2}=\varnothing$. Therefore, we obtain
@@
We finally show Eq. (S42). Since any $N\in M_{k}(e)$ is given by $N=N_{1}\cup N_{2}$ such that $N_{1}\in M_{k-1}(e)$ and $N_{2}\in M_{k-1}(e^{\prime})$, and $e$ and $e^{\prime}$ are $(d_{k-1}/2,b_{k-1}+d_{k-1}/2)$-linked, we obtain
@@
$|\displaystyle\absolutevalue{M_{k}(e)}$ $\leq\prod_{j=0}^{k-1}\left[\Lambda\quantity(b_{k-j-1}+\frac{d_{k-j-1}}{2})^{\Delta}\right]^{2^{j}}.$ (S50)